A rectangular garden has dimensions of 15 feet by 11 feet. A gravel path of uniform width is to be built around the garden. How wide can the path be if there is enough gravel for 192 square feet?
Answer by nerdybill(7384)

"area of outside edge of gravel path" - "area of garden" = 192
.
Let x = width of gravel path
then
(2x+15)(2x+11) - (15)(11) = 192
4x^2+22x+30x+165 - 165 = 192
4x^2+52x = 192
x^2+13x = 48
x^2+13x-48 = 0
factoring:
(x+16)(x-3) = 0
x = {-16, 3}
.
Tossing out the negative solution leaves us with:
width of path = 3 feet